3x+5y=15 and y=x+3
Answer by checkley77(12844)

y=x+3 now replace y in the secomd equation by (x+3) & solve for x.
3x+5y=15
3x+5(x+3)=15
3x+5x+15=15
8x=15-15
8x=0
x=0 ans
y=0+3
y=3 ans.
Proof:
3*0+5*3=15
15=15